OpenStudy (anonymous):

x - y = 6 x + y = 8 Solve system by graphing. I

OpenStudy (anonymous):

you can re arrange to have y= form x-y = 6 |+y x = 6+y |-6 x-6 = y -> y=x-6 "x + y = 8" subtract x from both sides -> y = 8-x

OpenStudy (anonymous):

the first equation has a positive slope (x is positive),the second equation has a negative slope (x is negative)

OpenStudy (anonymous):

AHHH. Thank you I just need to find some graphing paper, hahah